# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ 3.1. Remote Sensing/Machine Learning Taxonomy
Target Surveillance is a major focus in the literature. Though uncommon in satellite imagery due to the typical low resolution of the images, this trend might be explained by the proportion of airborne data. Indeed, Unmanned Aerial Vehicles (UAVs) enable better resolution and, subsequently, higher detection accuracy. To mitigate low accuracies, many experiments express Target Surveillance applications as scene-wise Classification, a solution relying on prudent patchification of the original full-size RS image888 RS images are too large for most ML models to process directly. The process of patchification divides them into smaller, more manageable patches.. Furthermore, we notice a significant focus on ”Cloud coverage detection.” Such application is motivated by the importance of clouds in optical images. Indeed, throughout the year, clouds cover over $66\%$ of Earth’s surface (Vitolo et al., 2024), significantly reducing the information obtainable from RGB images [^9] Conversely, SAR wavelengths are specifically chosen to be outside the absorption spectrum of clouds, allowing radar signals to pass through without significant attenuation.. Regarding ML tasks, most surveyed studies frame their applications as supervised Classification. In contrast, Regression appears only once, making it the most under-represented task. Similarly, most datasets used for Pixel classification are labeled and result in supervised problems, except for the ”Landcover/Land use” unsupervised classification of Gyaneshwar and Nidamanuri (2022). The Segmentation of urban areas follows the same trend, though Ratnakumar and Nanda (2021) use unsupervised clustering. Finally, all Object detection applications use multi-class datasets and single-shot models to detect multiple instances in a unique pass.
Landscape and environmental analyses are the prominent focus of the surveyed literature ($52\%$). Target surveillance, frequently simplified as scene-wise classification, dominates the rest of the research ($45\%$), with $30\%$ of the applications addressing defense concerns. Regarding data modalities, RGB data is dominant ($43\%$), with HSI remaining niche and SAR primarily used for water environments. 2.1 TAKEAWAYS
@@
[…]
[Task and ML model type distribution over the years]A stacked bar chart showing the distribution of different ML tasks and model types over time (from 2017 to 2024). The tasks are Classification, Pixel Classification, Segmentation, Object Detection, and Regression, each represented by a different color. The y-axis represents the number of experiments. Each bar represents a year, while each sub-bar corresponds to the models grouped per ”core”.
@@
Surprisingly, more applications are expressed as Pixel classification rather than Segmentation, despite the latter being the modern approach for pixel-level analysis (Asgari Taghanaki et al., 2021). Indeed, Segmentation networks, like U-Net (Ronneberger et al., 2015), have access to spatial information and generally outperform Pixel classification methods. However, resource constraints and computation overload also drive the model choice. Unlike deep Segmentation architectures, Pixel classification experiments often employ shallow networks or traditional ML, which are not only less demanding computationally but also easier to implement and parallelize. This combination of efficiency and simplicity likely explains their prevalence in this FPGA-focused survey, despite the broader RS trend favoring Segmentation. The rest of the problems are formulated as Classification and Object detection tasks. Together, these two categories represent over two-thirds of the experiments. Such a trend can be explained by the versatility of formulating problems as Classification tasks, as well as the importance of target monitoring in RS.
@@
Model cores are chosen to highlight the fundamental differences between architectures. In particular, traditional ML methods, such as SVMs (Cortes and Vapnik, 1995) or Decision Trees (Quinlan, 1986), are isolated from Neural Networks (NNs), just as Shallow Networks are distinguished from Deep Neural Networks (DNNs) (LeCun et al., 2015). We split CNNs—the vast majority of the models analyzed in this survey—between renowned architectures, such as LeNet (Lecun et al., 1998), AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2015), U-Net (Ronneberger et al., 2015), or ResNet (He et al., 2016), and custom CNN architectures. Notably, ”Custom CNN” includes both newly designed CNNs and named architectures used only once in this survey, such as SqueezeNet (Iandola et al., 2016) or ENet (Paszke et al., 2016). Finally, we classify YOLO (v2, v3, v4) (Redmon and Farhadi, 2017, 2018; Bochkovskiy et al., 2020) and SSD (Liu et al., 2016) architectures separately, even though renowned CNN architectures are frequently used as the backbones of detection networks. Indeed, the detection layers added to the feature extraction network (Fully Connected (FC) layers for YOLO and convolution layers for SSD) introduce fundamental differences to the architectures, such as the multiple heads of YOLO networks. Such characteristics reinforce the specialization of these models for object detection tasks.
@@ A CNN-Focused Landscape
Although shallow networks, such as MLPs, and traditional methods, like Decision Trees and SVMs, are featured in some studies (Gyaneshwar and Nidamanuri, 2022; Martins et al., 2024)—often due to their compatibility with resource-limited hardware—, the vast majority of contributions leverage DL algorithms. Within DL, convolutional architectures are strongly preferred: $76\%$ of all models are CNN-based. Models like AlexNet, ResNet, and VGG are occasionally used, often for benchmarking purposes, but custom-designed CNNs are the most prevalent ($27\%$ of all models), reflecting designers’ needs for medium-sized architectures typically with $6$ to $10$ layers. All detection networks are single-shot architectures like YOLO and SSD. Two-stage detection alternatives like R-CNN (Girshick et al., 2014) are not used in the analyzed experiments. Indeed, even if YOLO and SSD architectures are typically larger than R-CNN, the absence of a second pass through the networks highly reduces the computational complexity and processing time, making single-shot networks efficient solutions for resource-limited platforms like FPGAs (Zhao et al., 2023).
Selecting the Right Model
Model selection is influenced by several factors, including the intrinsic suitability of architectures for specific tasks (e.g., YOLO and SSD for object detection) and hardware constraints. For example, while SVMs can fit on most FPGA boards, larger models like ResNet50, with $\sim 25M$ parameters and high memory and computational demands, are unsuitable for smaller devices. This drives the prioritization of lightweight architectures, such as MobileNets (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019), and explains the large number of custom CNN designs (Li et al., 2019; Zhao et al., 2023). Lastly, the software used to integrate and deploy the method on the FPGA platform also drives the model choice. Indeed, if the designers are unfamiliar with low-level FPGA design, only the architectures already supported by higher-level toolchains are available. We further explore this aspect later, through  2.1 in Section 4.2.
With a steadily growing number of publications per year, formulating RS applications as Classification problems is prevailing ($36\%$). Of all the models, $76\%$ are convolution-based of which $27\%$ are custom CNNs with an average of $9$ layers. All detection models are single-shot networks and $85\%$ of the traditional ML methods are used in Pixel classification tasks. 2.1 TAKEAWAYS
@@
As explored in the previous sections, deploying AI-supported RS applications in resource-constrained environments requires formulating the application as a suitable ML task and selecting an appropriate model. These decisions are not only deeply intertwined with the selection of the hardware platform, but also driven by project requirements. For example, space deployment introduces harsh environmental conditions, including Single-Event Effects (SEEs) [^10] SEEs manifest as Single-Event Upsets (SEUs)—soft errors altering memory—or as Single-Event Latch-ups (SELs)—potentially destructive hard errors requiring a power reset. Refer to the supplemental material for developed space constraints. in onboard electronics caused by radiation exposure (Lange et al., 2024), tight power budgets from $20$W to $95$W for SmallSats payloads (Kothari et al., 2020), and extreme temperature swings from $-150^{\circ}$ to $+150^{\circ}$ (Gordo et al., 2020)—all of which place stringent demands on such edge computing hardware. In light of these challenges, this section explores FPGAs’ specific attributes and suitability for deploying AI on the edge compared to other hardware platform alternatives.
@@
[Bubble chart displaying FPGA boards used in article per year] A bubble chart showing the usage of different FPGA-based products over time. The x-axis represents the release year of the FPGA board, and the y-axis represents the number of DSPs available per device on a logarithmic scale. Each bubble represents a specific FPGA product, its color indicating the FPGA family (Zynq, Kria, Virtex, Alveo, Kintex, Artix, Spartan, or Cyclone). The size of each bubble is proportional to the number of times the board was used in the surveyed articles. The detailed FPGA part name is indicated under each bubble, and the legend shows the color-coding for each FPGA family. The bubbles follow a diagonal trend, indicating that newer boards tend to have more DSPs available.
Figure 3. Distribution of FPGAs used across articles. Unlike other figures/tables, it shows FPGA products used per article ($46$) instead of per experiment ($66$). * indicates additional boards used in the studies, absent from the experiments reported in Table 3.
@@
FPGAs are reconfigurable integrated circuits, that offer design flexibility at the cost of programming complexity. Often available off-the-shelf, FPGAs can operate independently on the edge or connect to other computational platforms via network cables or PCIe. Fig. 3 presents the different FPGAs used in each study. The devices are spread along the x-axis depending on their year of release, while the y-axis represents the number of DSP slices present on the board—relevant for arithmetic computations, like ML. Except for one experiment using the Cyclone V FPGA from Intel, we observe that all considered studies deploy hardware from AMD (formerly Xilinx). The most common family with $60\%$ are Zynq FPGAs, which package FPGA programmable logic with CPUs on a single die to support networking and general-purpose tasks. Such products will be referred to as System-on-Chip (SoC) FPGAs. In contrast, datacenter FPGAs, such as the Alveo Series, are often offered as PCIe cards that can be inserted into a server.
In addition to FPGAs, designers can select from several families of Processing Units (PUs) when deploying AI models on edge devices, each offering unique advantages. In the following subsections, we look at the alternatives: CPUs, GPUs, and ASICs (custom, TPUs, and VPUs). We analyze the authors’ motivations to select FPGA over alternatives and mention, when available, the comparisons made by the studies.
@@
Central Processing Units (CPUs) are ubiquitous general-purpose processing units, widely available as Commercial Off-The-Shelf (COTS) components. While well-suited for complex data transformations and system management, CPUs are limited in parallel processing capabilities, often falling behind specialized hardware in raw computational power. Evaluating FPGA-based designs against CPUs is a common practice in the surveyed studies. These comparisons often involve desktop and server processors, including older Intel i7 models (up to 8th gen) (Martins et al., 2024; Hashimoto et al., 2019; Li et al., 2023; Yang et al., 2023), newer i7/i9 models (12th and 13th gen) (Huang et al., 2024; Zhang et al., 2023a), AMD Ryzen 3990x (Zhang et al., 2022), and Broadwell Xeon server CPUs (Ni et al., 2023; Zhang et al., 2020; Wu et al., 2021). While these desktop and server CPUs offer high computational performance, they are not designed for energy-efficient applications, making direct comparisons with FPGA implementations less straightforward. In contrast, some studies include mobile CPUs, such as the AMD 4800H (Zhang et al., 2022; Zhao et al., 2023), which provide a more power-conscious alternative. Similarly, Huang et al. (2024) compare their ETAUS system against an embedded ARM Cortex A530 CPU and achieve significant performance improvements—×$\times 25$ GOP/s/W and ×$\times 37$ FPS.
Independently of direct comparison to CPUs, one of the most frequently cited advantages of FPGAs is their superior energy efficiency (Yang et al., 2023; Gyaneshwar and Nidamanuri, 2022; Yan et al., 2022; He et al., 2023; Li et al., 2023; Zhang et al., 2021; Chen et al., 2020b; Bahl et al., 2019; Li et al., 2020a; Matos-Carvalho et al., 2019). For instance, Yang et al. (2023) report a power efficiency ratio $29$ times superior to an Intel Core i7-8700 K. Indeed, FPGAs’ power consumption can be optimized using various design techniques, as detailed by Grover and K.Soni (2012). Furthermore, FPGAs are often favored for their radiation tolerance, with vendors like AMD offering radiation-hardened and radiation-tolerant Virtex and Kintex products. Although similar technologies are available for CPUs and microcontrollers, such as ARM Cortex-R or LEON families, FPGAs’ inherent reconfigurability facilitates the implementation of mitigation techniques, like Triple Modular Redundancy (TMR), where critical logic is triplicated and majority voting used to mask errors. On the downside, radiation-hardened FPGAs are significantly more expensive than rad-hard CPUs and microcontrollers. Their SRAM-based memory is also vulnerable to SEUs, requiring mitigation techniques (Sabogal and George, 2021). While investigating SRAM sensibility, Siegle et al. (2015) highlight that, although CPUs and microcontrollers remain essential onboard, especially for the OBC, modern payload processing applications increasingly rely on FPGAs. Over time, new solutions to deploy ML algorithms on embedded devices have been developed, shifting the focus from CPUs to domain-specific accelerators (Raoofy et al., 2021; Shawahna et al., 2019).
@@
Graphics Processing Units (GPUs) possess many small cores, enabling high parallelism and performance for tasks involving massive amounts of data and large matrix computations. They are widely used for training NNs. Consequently, many studies compare FPGA-based designs against high-performance NVIDIA GPUs. These include Maxwell (Wu et al., 2021; Hashimoto et al., 2019) and Pascal architectures (Ni et al., 2023; Zhang et al., 2020, 2023b; Yan et al., 2022; Zhang et al., 2021; Li et al., 2019), which introduce native half-precision support. Later architectures, such as Volta (Li et al., 2023; Myojin et al., 2020) and Turing (Nguyen et al., 2024; Zhao et al., 2023; Yang et al., 2023, 2022), contain Tensor Cores, processing units dedicated for Deep Learning operations. Ampere GPUs (Yu et al., 2024; Zhang et al., 2022, 2023a) introduce native support for low precision integer operations, thus heavily benefiting from quantization. The latest Blackwell GPUs (NVIDIA, 2025) support precision as low as 4-bit floating-point, decreasing FPGAs’ benefits gained from using custom datatypes. GPUs’ considerable processing capabilities come at the cost of significant memory and power consumption—often exceeding $300$W (Collange et al., 2009)—making them impractical for most edge deployments. To address these constraints, edge GPUs offer improved efficiency for AI inference. Computing boards, like the NVIDIA Jetson Nano (NVIDIA, 2019), integrate CPUs with energy-efficient GPUs microarchitectures. Yet, Huang et al. (2024) report that their FPGA-based system ETAUS achieves 2.7×2.7$\times$2.7 × GOP/s/W compared to Jetson Nano. Similarly, the AP2D-Net presented by Li et al. (2020a) achieves 2.8×2.8$\times$2.8 × FPS/W compared to the best model of the DAC System Design Contest 2018 (Xu et al., 2021) deployed on a Jetson TX2. Notably, Li et al. (2019) compare their FPGA implementations against the theoretical peak performance of edge GPUs like the Jetson AGX Xavier. They highlight that a significant gap in achievable performance still exists, even after extensive design optimization.
While GPUs deliver high computational throughput, FPGAs typically exhibit better power efficiency in edge scenarios (Wei et al., 2019). For instance, Li et al. (2023) report achieving $134\%$ of the FPS of a Tesla V100 while consuming 10×10$\times$10 × less power, whereas Li et al. (2019) find that an NVIDIA GTX 1070 Ti outperforms a Zynq-7000 in GOP/s/W efficiency by a factor of three. These variations highlight that power efficiency is highly implementation-dependent. The complex architecture and high transistor density of GPUs pose significant challenges for circuit-level radiation hardening. While the NVIDIA Tegra K1 (TK1) has demonstrated the ability to withstand radiation levels present in short-duration LEO missions (Badia et al., 2022). To the author’s knowledge, no edge GPU has been explicitly designed for space-grade reliability. The need for ionizing-radiation tolerance to ensure mission reliability is more readily addressed by the accessibility of radiation-tolerant FPGAs (Upadhyay et al., 2024). AMD’s KCU105 kit, for example, is a commercial equivalent of the space-grade Kintex Ultrascale XQRKU060 used in the VIDEO project—covered by Neris et al. (2022). In summary, while GPUs—especially edge GPUs—offer strong computational performance and power efficiency trade-offs for ML tasks, FPGAs remain the preferred choice in power- and space-constrained environments due to their energy efficiency and radiation tolerance.
@@
Application-Specific Integrated Circuits (ASICs) are fully custom-designed chips optimized for specific applications, potentially offering unmatched performance and energy efficiency. Yet, their design is a complex, lengthy, and expensive process—no study has compared an FPGA-based design against a suitable custom ASIC. Many studies list reconfigurability as a key advantage of FPGAs over ASICs (Ratnakumar and Nanda, 2021; Yang et al., 2023; Torres et al., 2020; Zhang et al., 2021; Li et al., 2020a; Yahiaoui et al., 2017); in defense and space for example, re-programmability enhances mission versatility by enabling the ”deploy and program” paradigm [^11] Allowing post-launch updates and in-orbit reconfiguration. (He et al., 2023; Li et al., 2023). Consequently, FPGAs are increasingly replacing traditional ASICs due to their faster time-to-market, cost-efficiency in small-batch applications, and competitive performance (Rapuano et al., 2021). Despite these advantages, ASICs remain the optimal choice for ultra-low-power applications, as the generic nature of FPGAs cannot match the efficiency of fully customized hardware (Amara et al., 2006).
Tensor Processing Units (TPUs) are commercially available ASICs optimized for Deep Learning, primarily used in data centers (Jouppi et al., 2017). While the Google Coral TPUs (Google, 2019) target edge devices, no study in this survey directly compared FPGA designs to a TPU. Li et al. (2019) only compare their Zynq 7000 design against Google cloud TPU’s theoretical peak performance. Vision Processing Units (VPUs), like Intel’s Movidius Myriad series, focus on low-power CNN inference. However, compared to the Myriad 2 VPU (Yang et al., 2022), FPGAs hold a key advantage in radiation tolerance. In a direct comparison, Rapuano et al. (2021) build on the CloudScout initiative (Giuffrida et al., 2020) and conclude that the Myriad 2 VPU is suitable only for short Low Earth Orbit (LEO) missions due to radiation susceptibility—unlike the space-grade Kintex Ultrascale XQRKU060 they also evaluate and is SEL-immune. Performance-wise, their design shows a 2.4×2.4$\times$2.4 × speed-up compared to the VPU, but with 1.8×1.8$\times$1.8 × higher power consumption and a longer development time.

### v2

@@ 3.1. Remote Sensing/Machine Learning Taxonomy
Target Surveillance is a major focus in the literature. Though uncommon in satellite imagery due to the typical low resolution of the images, this trend might be explained by the proportion of airborne data. Indeed, Unmanned Aerial Vehicles (UAVs) offer higher-resolution imagery, which supports higher detection accuracy. To mitigate low accuracies, many experiments express Target Surveillance applications as scene-wise Classification, a solution relying on prudent patchification of the original full-size RS image [^9] RS images are too large for most ML models to process directly. The process of patchification divides them into smaller, more manageable patches.. Furthermore, we notice a significant focus on estimating cloud coverage. Such an application is motivated by the importance of clouds in optical images. Indeed, throughout the year, clouds cover over $66\%$ of Earth’s surface (Vitolo et al., 2024), significantly reducing the information obtainable from RGB images [^10] Conversely, SAR wavelengths are specifically chosen to be outside the absorption spectrum of clouds, allowing radar signals to pass through without significant attenuation.. Regarding ML tasks, most surveyed studies frame their applications as supervised Classification. In contrast, Regression appears only twice, making it the most under-represented task, although retrieving parameters from a scene is a critical problem in RS. Similarly, most datasets used for Segmentation - Pixel are labeled and result in supervised problems, except for the LULC unsupervised classification of Gyaneshwar and Nidamanuri (2022). The semantic segmentation of urban areas follows the same trend, though Ratnakumar and Nanda (2021) use unsupervised clustering. Finally, all Object detection applications use multi-class datasets and single-shot models to detect multiple instances in a unique pass.
Fully updated. Target surveillance is the prominent focus of the surveyed literature ($44\%$). Frequently simplified as scene-wise classification, $31\%$ of the applications address defense concerns. Landscape and environment monitoring dominates the rest of the research ($40\%$). Regarding data modalities, RGB data is dominant ($65\%$), with HSI remaining niche and SAR primarily used for water environments. 2.1 TAKEAWAYS
@@
Refer to caption Task and ML model type distribution over the years A stacked bar chart showing the distribution of different ML tasks and model types over time (x-axis from 2017 to 2024). The tasks are Classification, Pixel Classification, Segmentation, Object Detection, and Regression, each represented by a different color. The y-axis represents the number of experiments. Each bar represents a year, while each sub-bar corresponds to the models grouped per ”core”.
@@
Surprisingly, among semantic segmentation tasks, more models operate per-pixel rather than processing the input patch at once, despite the latter being the modern approach for pixel-level analysis (Asgari Taghanaki et al., 2021). Indeed, segmentation networks, like U-Net (Ronneberger et al., 2015), have access to spatial information and generally outperform methods limiting spatial neighborhood information. However, resource constraints and computation overload also drive the model choice. Unlike Segmentation - Patch deep architectures, Segmentation - Pixel experiments often employ shallow networks or traditional ML, which are not only less demanding computationally but also easier to implement and parallelize. This combination of efficiency and simplicity likely explains their prevalence in this FPGA-focused survey, despite the broader RS trend favoring Segmentation - Patch. The remaining problems are formulated as Classification and Object detection tasks. Together, these two categories represent two-thirds of the experiments. Such a trend can be explained by the versatility of formulating problems as Classification tasks, as well as the importance of target monitoring in RS.
@@
Model cores are chosen to highlight the fundamental differences between architectures. In particular, traditional ML methods, such as SVMs (Cortes and Vapnik, 1995) or Decision Trees (Quinlan, 1986), are isolated from Neural Networks (NNs), just as Shallow Networks are distinguished from Deep Neural Networks (DNNs) (LeCun et al., 2015). We split CNNs—the vast majority of the models analyzed in this survey—between renowned architectures, such as LeNet (Lecun et al., 1998), AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2015), U-Net (Ronneberger et al., 2015), or ResNet (He et al., 2016), and custom CNN architectures. Notably, ”Custom CNN” includes both newly designed CNNs and named architectures used only once in this survey, such as SqueezeNet (Iandola et al., 2016) or ENet (Paszke et al., 2016). Finally, we classify YOLO (v2, v3, v4) (Redmon and Farhadi, 2017, 2018; Bochkovskiy et al., 2020) and SSD (Liu et al., 2016) architectures separately, even though renowned CNN architectures are frequently used as the backbones of detection networks. Indeed, the detection layers added to the feature extraction network (Fully Connected (FC) layers for YOLO and convolution layers for SSD) introduce fundamental differences to the architectures, such as the multiple heads of YOLO networks.
@@ A CNN-Focused Landscape
Although shallow networks, such as MLPs, and traditional methods, like Decision Trees and SVMs, are featured in some studies (Gyaneshwar and Nidamanuri, 2022; Martins et al., 2024)—often due to their compatibility with resource-limited hardware—, the vast majority of contributions leverage DL algorithms. Within DL, convolutional architectures are strongly preferred. Models like AlexNet, ResNet, and VGG are occasionally used, often for benchmarking purposes, but custom-designed CNNs are the most prevalent ($27\%$ of all models), reflecting designers’ needs for medium-sized architectures typically with $6$ to $10$ layers. All detection networks are single-shot architectures like YOLO and SSD. Two-stage detection alternatives like R-CNN (Girshick et al., 2014) are not used in the analyzed experiments. Indeed, even if YOLO and SSD architectures are typically larger than R-CNN, the absence of a second pass through the networks highly reduces the computational complexity and processing time, making single-shot networks efficient solutions for resource-limited platforms like FPGAs (Zhao et al., 2023).
With a steadily growing number of publications per year, formulating RS applications as Classification problems is prevailing ($35\%$). Of all the models, $74\%$ are convolution-based, of which $26\%$ are custom CNNs with an average of 99 layers. All detection models are single-shot networks and $89\%$ of the traditional ML methods are used in Segmentation - Pixel tasks. 2.1 TAKEAWAYS
@@
As explored in the previous sections, deploying AI-supported RS applications in resource-constrained environments requires formulating the application as a suitable ML task and selecting an appropriate model. These decisions are deeply intertwined with the selection of the hardware computing platform.
Hardware Constraints Across RS Platforms
RS platforms are vehicles that carry remote sensors; they can operate at various altitudes as spaceborne, airborne, or ground-based systems. Each platform operates under distinct hardware constraints. Size, Weight, Power, and Cost (SWaP-C) define the fundamental trade-offs between computational capability and mission feasibility. Additionally, payloads may require radiation hardening or thermal management. For example, spaceborne systems face the most stringent constraints. They are exposed to harsh environmental conditions, including Single-Event Effects (SEEs) [^11] SEEs manifest as Single-Event Upsets (SEUs)—soft errors altering memory—or as Single-Event Latch-ups (SELs)—potentially destructive hard errors requiring a power reset. in onboard electronics caused by radiation exposure (Lange et al., 2024), extreme temperature swings from $150^{\circ}$C to $150^{\circ}$C (Gordo et al., 2020), and tight power budgets from 2020W to 9595W for SmallSats payloads (Kothari et al., 2020). These extreme conditions make spaceborne systems the most constrained environment for FPGA-based ML deployment and, accordingly, the primary focus of the surveyed studies. Airborne platforms, targeted by $\sim\!25\%$ of the articles, present moderate constraints that differ significantly between sub-categories (Watts et al., 2012). Micro and nano UAVs operate at low altitudes and have short endurance, typically requiring light payloads and possibly high power if taking off vertically. On the other hand, high- and medium-altitude aircraft can typically accommodate substantial payloads and possess longer endurance. Finally, ground-based systems operate with minimal SWaP-C constraints and are absent from this review. Given these platform-specific limitations, this section examines how FPGAs address such challenges and how they compare with conventional computing hardware for edge AI deployment.
@@
Figure 4. Distribution of FPGAs used across articles. Unlike other figures/tables, it shows FPGA products used per article ($48$) instead of per experiment ($68$). * indicates additional boards used in the studies, absent from the experiments reported in Table 4. The horizontal dotted lines represent arbitrary distinctions between size categories used in RQ7.
@@
FPGAs are reconfigurable integrated circuits, that offer design flexibility at the cost of programming complexity. Often available off-the-shelf, FPGAs can operate independently on the edge or connect to other computational platforms via network cables or PCIe. Fig. 4 presents the different FPGAs used in each study. The devices are spread along the x-axis depending on their year of release, while the y-axis represents the number of DSP slices present on the board—relevant for arithmetic computations, like ML. Except for one experiment using the Cyclone V FPGA from Intel, we observe that all considered studies deploy hardware from AMD (formerly Xilinx). The most common family are Zynq FPGAs, used by $57\%$ of the studies, which package FPGA programmable logic with CPUs on a single die to support networking and general-purpose tasks. Such products will be referred to as System-on-Chip (SoC) FPGAs. In contrast, datacenter FPGAs, such as the Alveo Series, are often offered as PCIe cards that can be inserted into a server. Furthermore, to address the need for ionizing-radiation tolerance to ensure mission reliability, FPGA vendors offer radiation-hardened and radiation-tolerant variants. AMD’s KCU105 kit, for example, is a commercial equivalent of the space-grade Kintex Ultrascale XQRKU060 used in the VIDEO project (Neris et al., 2022). In addition to FPGAs, designers can select from several families of Processing Units (PUs) when deploying AI models on edge devices. In the following subsections, we look at the alternatives: CPUs, GPUs, and ASICs (custom, TPUs, and VPUs). We analyze the authors’ motivations to select FPGA over alternatives and mention, when available, the comparisons made by the studies. In the end, we summarize the unique advantages of each family of PUs in Table 3.
@@
[…]
@@
Graphics Processing Units (GPUs) possess many small cores, enabling high parallelism and performance for tasks involving massive amounts of data and large matrix computations. They are widely used for training NNs. Consequently, many studies compare FPGA-based designs against high-performance NVIDIA GPUs. These include Maxwell (Wu et al., 2021; Hashimoto et al., 2019) and Pascal architectures (Ni et al., 2023; Zhang et al., 2020, 2023; Yan et al., 2022; Zhang et al., 2021; Li et al., 2019b), which introduce native half-precision support. Later architectures, such as Volta (Li et al., 2023a; Myojin et al., 2020) and Turing (Nguyen et al., 2024; Zhao et al., 2023; Yang et al., 2023, 2022), contain Tensor Cores, processing units dedicated for Deep Learning operations. Ampere GPUs (Yu et al., 2024; Zhang et al., 2022, 2023) introduce native support for low precision integer operations, thus heavily benefiting from quantization. The latest Blackwell GPUs (NVIDIA, 2025) support precision as low as 4-bit floating-point, decreasing FPGAs’ benefits gained from using custom datatypes. FPGAs can exhibit better power efficiency (Wei et al., 2019) compared to desktop GPUs. For example, Li et al. (2023a) report $134\%$ of the FPS of a Tesla V100 while consuming $10\times$ less power. In contrast, Li et al. (2019b) find that an NVIDIA GTX 1070 Ti outperforms a Zynq-7000 in GOP/s/W efficiency by a factor of three showing, that power efficiency is highly implementation-dependent. Regardless of efficiency, GPUs’ considerable processing capabilities come at the cost of significant memory and power consumption—often exceeding 300300W (Collange et al., 2009)—making them impractical for most edge deployments. To address these constraints, computing boards, like the NVIDIA Jetson Nano (NVIDIA, 2019), combine CPUs with energy-efficient edge GPUs. Even though edge GPUs usually deliver higher peak performance (Li et al., 2019b), FPGAs can still achieve better power efficiency (Wei et al., 2019). For example, Huang et al. (2024) report that their FPGA-based system ETAUS achieves $2.7\times$ GOP/s/W compared to Jetson Nano. Similarly, the AP2D-Net presented by Li et al. (2020b) achieves $2.8\times$ FPS/W compared to the best model of the DAC System Design Contest 2018 (Xu et al., 2021) deployed on a Jetson TX2. The complex architecture and high transistor density of GPUs pose significant challenges for circuit-level radiation hardening. While the NVIDIA Tegra K1 has withstood radiation levels present in short-duration LEO missions (Badia et al., 2022), no edge GPU, to the authors’ knowledge, has been explicitly designed for space-grade reliability. In summary, while GPUs—especially edge GPUs—offer strong computational performance and power efficiency trade-offs for ML tasks, FPGAs remain the preferred choice in power- and space-constrained environments due to their energy efficiency and radiation tolerance.
@@
Application-Specific Integrated Circuits (ASICs) are fully custom-designed chips optimized for specific applications, potentially offering unmatched performance and energy efficiency. Yet, their design is a complex, lengthy, and expensive process—no study has compared an FPGA-based design against a suitable custom ASIC. Many studies list reconfigurability as a key advantage of FPGAs over ASICs (Ratnakumar and Nanda, 2021; Yang et al., 2023; Torres et al., 2020; Zhang et al., 2021; Li et al., 2020b; Yahiaoui et al., 2017); in defense and space for example, re-programmability enhances mission versatility by enabling the ”deploy and program” paradigm [^12] Allowing post-launch updates and in-orbit reconfiguration. (He et al., 2023; Li et al., 2023a). Consequently, FPGAs are increasingly replacing traditional ASICs due to their faster time-to-market, cost-efficiency in small-batch applications, and competitive performance (Rapuano et al., 2021). Despite these advantages, ASICs remain the optimal choice for ultra-low-power applications, as the generic nature of FPGAs cannot match the efficiency of fully customized hardware (Amara et al., 2006). Tensor Processing Units (TPUs) are commercially available ASICs optimized for Deep Learning, primarily used in data centers (Jouppi et al., 2023). While the Google Coral TPUs (Google, 2019) target edge devices, no study in this survey directly compared FPGA designs to a TPU. Li et al. (2019b) only compare their Zynq 7000 design against Google cloud TPU’s theoretical peak performance. Vision Processing Units (VPUs), like Intel’s Movidius Myriad series, focus on low-power CNN inference. However, compared to the Myriad 2 VPU (Yang et al., 2022), FPGAs hold a key advantage in radiation tolerance. In a direct comparison, Rapuano et al. (2021) build on the CloudScout initiative (Giuffrida et al., 2020) and conclude that the Myriad 2 VPU is suitable only for short Low Earth Orbit (LEO) missions due to radiation susceptibility—unlike the space-grade Kintex Ultrascale XQRKU060 they also evaluate and is SEL-immune. Performance-wise, their design shows a $2.4\times$ speed-up compared to the VPU, but with $1.8\times$ higher power consumption and a longer development time.
Table 3. Summary of different computing platforms used in the different studies.
Chip Dev. Time Price/Unit Rad-Hard (Avail.) Updatable Pow. Eff. Latency Peak Perf. (fp32)
CPU Short Low Yes Yes Low Medium Medium
GPU Medium Medium/High No Yes Medium High Very high
Custom ASIC Very long Very low Yes No Very High Very Low Custom
FPGA Long High Yes Yes High Low Medium